The sum of the digits of a two digit number is 14. The difference between the tens digit and the units digit is 2. What is the number?
Answer by opeodu(24)

let the ten digit be 'X' and the unit digit be 'Y'.
X+Y=14-----(A)
X-Y=2------(B)
2X=16
X=16/2=8
Y=14-X
Y=14-8=6